OpenStudy (anonymous):

Find the surface area of an orange with a 2'' radius.

OpenStudy (anonymous):

I AM DOING 4 X 22/7 X 4 AND IT KEEPS SHOWING THE WRONG ANSWER

OpenStudy (apoorvk):

well, what is your final calculated answer?

OpenStudy (anonymous):

352/7

OpenStudy (anonymous):

SA = 4πr^2 SA = 4(22/7)(2)^2 SA = 4(22/7)(4) SA = 16(22/7) SA = 352/7 You just have to plug the numbers into the formula.

OpenStudy (anonymous):

WHEN I PUT 352/7 IT IS SAYING ITS THE WRONG ANSWER

OpenStudy (apoorvk):

Your answer seems correct. some error on their part I guess.

OpenStudy (anonymous):

Try putting it in as a decimal? 352/7 ≈ 50.2857

OpenStudy (anonymous):

Do you need a more precise value of pi? Also, are you assuming the orange is perfectly smooth, or do you have to consider the texture?

OpenStudy (anonymous):

LET ME TRY 50.2857 AND SEE NOPE ITS SAYING ITS WRONG

OpenStudy (anonymous):

Using a more precise value of pi, it is closer to 50.265. Do you know how many decimal digits of precision you need to report, or should you leave it as an exact value (in terms of pi)?

OpenStudy (anonymous):

its not telling me

OpenStudy (anonymous):

i am also having trouble with this one as well Two concentric spheres have radii of 5" and 6". Find the volume of the space between them. cubic inches

OpenStudy (anonymous):

Take the volume of the bigger one and subtract the volume of the smaller one. V = 4/3πr^3

OpenStudy (anonymous):

4/3 pie 1 = 4/3 right

OpenStudy (anonymous):

No. You can't just subtract the radii unfortunately...

OpenStudy (anonymous):

Take the volume of the bigger sphere: V = 4/3πr^3 V = 4/3π(6)^3 V = 288π Then find the volume of the smaller sphere. V = 4/3πr^3 V = 4/3π(5)^3 V = 500π/3 Subtract the two volumes to get what we call an annulus. 288π - 500π/3 = 364π/3 or if you want an approximate answer, just use 3.14 or 22/7 for π. Btw, π is spelled pi not pie lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yes thanks my lesson just told me the answer for the previous question it is saying the answer is 16

OpenStudy (anonymous):

not sure how they figured that one out.

OpenStudy (anonymous):

That means to leave it in terms of pi. The SA=16pi.

OpenStudy (anonymous):

That would be the surface area without the π. lol. Probably just a coincidence.

OpenStudy (anonymous):

Maybe not.

OpenStudy (anonymous):

oh okay i just figured it out wrong

OpenStudy (anonymous):

Yeah. Just a clerical error.